OpenStudy (anonymous):

State the property that justifies the following: 9(x + 3) = 9x + 27

OpenStudy (anonymous):

Additive Inverse Distributive Property Associate Property of Addition Commutative Property of Addition

OpenStudy (anonymous):

when you multiply an outer number (in your case 9) by the values in the parentheses, do you know what you are doing?

OpenStudy (anonymous):

distributing

OpenStudy (anonymous):

The distributive property because you are distributing the 9 to the x and the 3

OpenStudy (anonymous):

perfect, so which property do you think it is?

OpenStudy (anonymous):

so its distributive

OpenStudy (anonymous):

if youre distributing, then it kind of already gives you your answer

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

what about this one

OpenStudy (anonymous):

3)(6)(9) = (6)(9)(3) Multiplicative Identity Additive Inverse Distributive Property Associate Property of Addition Commutative Property of Addition Commutative Property of Multiplication Associate Property of Multiplication

OpenStudy (anonymous):

communitive property of multiplication

OpenStudy (anonymous):

cool thought so

OpenStudy (anonymous):

Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4)

OpenStudy (anonymous):

i disagree with that @sammyx335

OpenStudy (anonymous):

ohh its associative

OpenStudy (anonymous):

yea, thats what i would put

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about this

ganeshie8 (ganeshie8):

careful, you're changing the positions of factors

ganeshie8 (ganeshie8):

(3)(6)(9) = (6)(9)(3) ab = bc is what property ? :)

OpenStudy (anonymous):

im having trouble understanding what you mean @ganeshie8

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

in this example, there are three different variables that are being moved in their position of multiplication. in that, would it not be associative?

ganeshie8 (ganeshie8):

associative is about grouping

OpenStudy (anonymous):

Where is Lord Ganesh?

ganeshie8 (ganeshie8):

a(bc) = (ab)c is associative property

ganeshie8 (ganeshie8):

Notice that we're not changing the position of numbers here

ganeshie8 (ganeshie8):

variables*

OpenStudy (anonymous):

so thats what this is(9)[(8)(7)] = [(9)(8)](7)

OpenStudy (anonymous):

im still confused because (a)(bc) will still get you the same thing if a=1, b=2, c=3, 1(2*3) = (1*2)(3)

OpenStudy (aum):

3 x 6 x 9 = (3x6) x 9 = 3 x (6x9) (associative) 3 x (6x9) = (6x9) x 3 (commutative)

ganeshie8 (ganeshie8):

Yes alygirl, (9)[(8)(7)] = [(9)(8)](7) is associative property

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

yes @TechnoSoul , thats the reason they are called properties of `equalities` :)

OpenStudy (anonymous):

haha i understand that part, I'm just still a bit confused. oh well, its alright. I dont want to confuse @alyygirl anymore than i have already XD